When the subjects were stratified by sex, the following tables were obtained:

Females:

Cases

Controls

      Sedentary lifestyle

15

30

      Not sedentary lifestyle

30

120

45

150

Males:

Cases

Controls

      Sedentary lifestyle

120

120

      Not sedentary lifestyle

15

30

135

150

(d) Does the relationship between sedentary lifestyle and heart disease appear to differ for females and males? (That is, is there an interaction between sex, sedentary lifestyle, and heart disease?) Why or why not?
Answer #1

solution:

Given data:

d)

. risk factor for females. 15/45 =1/3

risk factor for males is 120/150 =4/3

Hence it is different for males and females., males are more prone .hence sex is a confounding factor.

please give me thumb up